OpenStudy (anonymous):

Can someone hel me please ? Solve the system of equation x+2y=8 4y=-2x+16

OpenStudy (phoenixfire):

Rearrange the first equation for x. Substitute into the second and solve for x.

OpenStudy (phoenixfire):

solve for y I mean.

OpenStudy (anonymous):

and what i do with the 4 in front of the y?

OpenStudy (phoenixfire):

x+2y=8 rearrange y=(8-x)/2 sub 4y=-2x+16 4(8-x)=-2x+16 solve for x 16-4x=-2x+16 16-16=-2x+4x 0=2x 0=x sub x back into a equation to find y x+2y=8 0+2y=8 y=8/2=4 x=0 y=4 sub both into other equation to verify 4(4)=-2(0)+16 16=16

OpenStudy (anonymous):

thanks so much

OpenStudy (phoenixfire):

No problem.

OpenStudy (anonymous):

systems equation can be solved by either elimination method or substitution method